OpenStudy (jchick):

Choose the equation below that represents the line passing through the point (1, -4) with a slope of one half. Choose one answer. a. y - 4 = one half(x - 1) b. y + 4 = 2(x - 1) c. y + 4 = one half(x - 1) d. y - 4 = 2(x + 1)

OpenStudy (jchick):

I choose B correct @Loser66

OpenStudy (jchick):

@Love_Ranaa am I correct?

OpenStudy (jchick):

@Nerd94 can you help me am I correct?

OpenStudy (anonymous):

I believe u may be right

OpenStudy (okra77):

no the correct answer is c

OpenStudy (jchick):

Are you sure?

OpenStudy (jchick):

How did you get that answer?

OpenStudy (jchick):

@okra77 how did you get that answer?

OpenStudy (okra77):

the equation is y-y1 = m( x-x1) so if the slope is 1/2 just sub in with the point and the slope so y+4=1/2(x-1)

OpenStudy (jchick):

Ok that makes sense thanks!

OpenStudy (okra77):

no problem

OpenStudy (jchick):

Will you help me with another one I mentioned you in it?